OpenStudy (anonymous):

write a problem you can represent with 2k-10=6

OpenStudy (anonymous):

do you already know it

OpenStudy (anonymous):

2k-10 = 6 Add 10 to both sides of equation 2k = 16 Divide both sides by 2 k = 8 To check the answer, substitute it in to the original equation: 2(8)-10 = 6 16-10 = 6 6 = 6

OpenStudy (anonymous):

@donttakemyheadphones r u there